A generic salt, AB3, has a molar mass of 185 g/mol and a solubility of 5.50 g/L at 25 °C.

Part A)

A generic salt, AB3, has a molar mass of 185 g/mol and a solubility of 5.50 g/L at 25 °C. 

AB3(s) ⇌ A3+ (aq) + 3B-(aq) 

What is the Ksp of this salt at 25 °C?


Part B)

A generic salt, AB2, has a molar mass of 271 g/mol and a solubility of 2.30 g/L at 25 °C. 

AB2(s) ⇌ A2+(aq) + 2B- (aq) 

What is the Ksp of this salt at 25 °C? 

Part C)

At 25 °C only 0.0890 mol of the generic salt AB3 is soluble in 1.00 L of water. 

AB3(s) ⇌ A3+ (aq) +3B- (aq) 

What is the Ksp of the salt at 25 °C? 
Answer #1

A)

Molar mass of AB3= 185 g/mol
s = 5.5 g/L
To covert it to mol/L, divide it by molar mass
s = 5.5 g/L / 185 g/mol
s = 2.973*10^-2 mol/L

At equilibrium:
AB3       <---->     A3+      +         3 B-    

                     s                  3s      


Ksp = [A3+][B-]^3
Ksp = (s)*(3s)^3
Ksp = 27(s)^4
Ksp = 27(2.973*10^-2)^4
Ksp = 2.109*10^-5
Answer: 2.11*10^-5

B)

Molar mass of AB2= 271 g/mol
s = 2.3 g/L
To covert it to mol/L, divide it by molar mass
s = 2.3 g/L / 271 g/mol
s = 8.487*10^-3 mol/L

At equilibrium:
AB2       <---->     A2+      +         2 B-    

                     s                  2s      


Ksp = [A2+][B-]^2
Ksp = (s)*(2s)^2
Ksp = 4(s)^3
Ksp = 4(8.487*10^-3)^3
Ksp = 2.445*10^-6
Answer: 2.44*10^-6

C)

At equilibrium:
AB3       <---->     A3+      +         3 B-    

                     s                  3s      


Ksp = [A3+][B-]^3
Ksp = (s)*(3s)^3
Ksp = 27(s)^4
Ksp = 27(8.9*10^-2)^4
Ksp = 1.694*10^-3
Answer: 1.69*10^-3